# PreP-OCR: A Complete Pipeline for Document Image Restoration and Enhanced OCR Accuracy

### v1

@@ Abstract
This paper introduces PreP-OCR, a two-stage pipeline that combines document image restoration with semantic-aware post-OCR correction to improve text extraction from degraded historical documents. Our key innovation lies in jointly optimizing image clarity and linguistic consistency. First, we generate synthetic image pairs with randomized text fonts, layouts, and degradations. An image restoration model is trained on this synthetic data, using multi-directional patch extraction and fusion to process large images. Second, a ByT5 post-corrector, fine-tuned on synthetic historical text training pairs, addresses any remaining OCR errors. Detailed experiments on 13,831 pages of real historical documents in English, French, and Spanish show that PreP-OCR pipeline reduces character error rates by 63.9–70.3% compared to OCR on raw images. Our pipeline demonstrates the potential of integrating image restoration with linguistic error correction for digitizing historical archives. https://github.com/NikoGuan/PreP-OCR
@@ 6 Conclusion
In this paper we proposed PreP-OCR, a synthetic-data-driven pipeline that restores images and improves text extraction from degraded historical documents. A key component of this work is the introduction of a synthetic data generation method that simulates realistic document degradations and typographic variations. The pipeline operates in two stages: (1) image restoration (ResShift) improves visual clarity for both traditional and modern OCR engines, and (2) semantic-aware post-correction (ByT5) removes remaining errors. Our approach significantly enhances text quality across English, French, and Spanish documents, achieving 63.9–70.3% CER reduction compared to raw OCR outputs.

### v2

@@ Abstract
This paper introduces PreP-OCR, a two-stage pipeline that combines document image restoration with semantic-aware post-OCR correction to enhance both visual clarity and textual consistency, thereby improving text extraction from degraded historical documents. First, we synthesize document-image pairs from plaintext, rendering them with diverse fonts and layouts and then applying a randomly ordered set of degradation operations. An image restoration model is trained on this synthetic data, using multi-directional patch extraction and fusion to process large images. Second, a ByT5 post-OCR model, fine-tuned on synthetic historical text pairs, addresses remaining OCR errors. Detailed experiments on 13,831 pages of real historical documents in English, French, and Spanish show that the PreP-OCR pipeline reduces character error rates by 63.9-70.3% compared to OCR on raw images. Our pipeline demonstrates the potential of integrating image restoration with linguistic error correction for digitizing historical archives. https://github.com/NikoGuan/PreP-OCR
@@ 6 Conclusion
In this paper we proposed PreP-OCR, a synthetic-data-driven pipeline that restores images and improves text extraction from degraded historical documents. A key component of this work is the introduction of a synthetic data generation method that simulates realistic document degradations and typographic variations. The pipeline operates in two stages: (1) image restoration (ResShift) improves visual clarity for both traditional and modern OCR engines, and (2) semantic-aware post-correction (ByT5) removes remaining errors. Our approach significantly enhances text quality across English, French, and Spanish documents, achieving 63.9-70.3% CER reduction compared to raw OCR outputs.